Find the prime factorization of the integers 1234, 10140, and 36000


1
Expert's answer
2021-09-21T02:17:57-0400

Consider the following integers:

1234,10140, and 36000


The objective is to find the prime factorization of the above-mentioned integers.

Recall that a positive integer can be uniquely expressed as the product of prime numbers.


The prime factorization of the given number is as follows:

"1234= 2\times617 \\

10140= 2\times2\times3\times5\times13\times13=2^2\times3\times5\times13^2 \\

36000=2\times2\times2\times2\times2\times3\times3\times5\times5\times5=2^5\times3^2\times5^3"


Therefore, the prime factorization of the given integers are

"1234= 2\times617 \\

10140=2^2\times3\times5\times13^2 \\

36000=2^5\times3^2\times5^3"